OpenStudy (anonymous):

4(xy-1)-(x-3) Simplify.

OpenStudy (unklerhaukus):

distribute the four across the first brackets, and distribute the - across the second brackets, then combine like terms

OpenStudy (anonymous):

I have no clue

OpenStudy (unklerhaukus):

\[4(xy-1)-(x-3)=(4\times xy+4\times-1)+(-1\times x+-1\times-3)\]